# Flow-Attentional Graph Neural Networks

### v2

@@ 1 Introduction
Graph neural networks (GNNs) (Scarselli et al., 2008; Kipf & Welling, 2017) have emerged as a powerful framework that extends the scope of deep learning from Euclidean to graph-structured data, which is prevalent across many real-world domains, such as social networks (Fan et al., 2019), recommender systems (Wu et al., 2022), materials science (Reiser et al., 2022), or epidemiology (Liu et al., 2024). Especially attention-based GNNs have become increasingly popular due to their ability to select relevant features adaptively (Sun et al., 2023). As graph data becomes increasingly common, advancing GNN architectures is crucial for improving performance in tasks such as node classification (Hamilton et al., 2017), graph regression (Gilmer et al., 2017), or link prediction (Zhang & Chen, 2018).
In many important applications of GNNs, graphs are naturally associated with a flow of physical resources, such as electrical current in electronic circuits (Sánchez et al., 2023) or power grids (Liao et al., 2021), traffic in transportation networks (Jiang & Luo, 2022), water in river networks (Sun et al., 2021), or raw materials and goods in supply chains (Kosasih & Brintrup, 2022). All nodes in these resource flow graphs, except for source and target nodes, are subject to Kirchhoff’s first law, which states that the sum of all incoming and outgoing flows must be zero, reflecting the conservation of resources. By contrast, informational graphs—such as computation graphs, social networks, or citation networks—are not associated with any physical flow but rather represent relationships or information transfer. Information can be arbitrarily duplicated in these graphs, unlike in flow graphs, where such duplication would violate the conservation law.
@@
Figure 1: Two non-isomorphic graphs that are equivalent as informational graphs, but different as resource flow graphs. a The two different directed graph structures represent the same computation (example adapted from Zhang et al. (2019)). b The same graph structures as above represent different electronic circuits.
@@ 2 Related Work
In recent years, many new GNN models have been specifically designed for different graph types (Thomas et al., 2023). Despite their fundamental differences, informational graphs and flow graphs are mainly treated by the same basic message-passing layers, such as GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), or GIN (Xu et al., 2019). In these models, messages exchanged between neighboring nodes do not depend on the number of recipients. Instead, the information is arbitrarily duplicated and passed to all neighbors. GCN (Kipf & Welling, 2017) applies a symmetric neighborhood normalization but exhibits limited expressivity and performance. Attention-based GNNs, such as GAT, GATv2 (Brody et al., 2022), or UniMP (Shi et al., 2021), adaptively weight neighboring node features, leading to improved representation learning. However, the attention weights are obtained through normalization across incoming neighbors, allowing for arbitrary message duplication. Our approach normalizes across outgoing neighbors instead, which better captures the physical flow of a graph while preserving the benefits of graph attention.
@@
Many flow graphs, including the example graphs in Fig. 1, can be naturally expressed as DAGs, e.g., operational amplifiers (Dong et al., 2023) or material flow networks (Perera et al., 2018). Before GNNs were introduced, recursive neural networks were applied to DAGs (Sperduti & Starita, 1997; Frasconi et al., 1998), and contextual recursive cascade correlation was proposed to overcome limitations in expressivity (Hammer et al., 2005). However, these early works lacked the advantages of modern GNNs, which have been extended to DAGs in recent years (Zhang et al., 2019; Thost & Chen, 2021). In directed acyclic GNNs, nodes are typically updated sequentially following the partial order of the DAG, and the final target node representations are used as the graph embedding. Although these sequential models outperform undirected GNNs on DAG datasets, they still aggregate node neighborhood information similarly. This means they only encode the computation tree of the output nodes and not the exact structure of the DAG, resulting in limited expressivity.
@@
A possible approach to overcome the problem of indistinguishable flow graphs is to use node indices or random features as input node features (Loukas, 2020; Sato et al., 2021), which enables the model to uniquely identify each node. However, the resulting GNN model is no longer permutation invariant, which reduces its generalization capability. Similar problems arise for Transformer-based models (Vaswani et al., 2017) such as PACE (Dong et al., 2022), which incorporate the relational inductive bias (Battaglia et al., 2018) via positional encodings. A different strategy would be to introduce Kirchhoff’s first law through an additional physics-informed loss term (Donon et al., 2020), which considerably increases the training complexity and is only useful if the target variable is the resource flow itself. Our approach enhances the expressivity of attention-based GNNs on flow graphs while preserving permutation invariance and computational efficiency.
@@ 3 Preliminaries
Node Multiset. For each node in a graph, the feature vectors of a set of incoming nodes can be represented as a multiset (Xu et al., 2019). A multiset is a pair $(S,m)$, where $S$ is a set of distinct elements (the node features) and $m:S\rightarrow\mathbb{N}$ is the multiplicity of each element. We call two multisets $X_{1}=(S,m_{1}),~X_{2}=(S,m_{2})$ equally distributed if $m_{2}=k\cdot m_{1}$ with $k\in\mathbb{N}_{\geq 1}$.
@@
The most widely used directed acyclic GNNs are D-VAE (Zhang et al., 2019) and DAGNN (Thost & Chen, 2021), the latter of which uses standard attention for aggregation. Both models utilize gated recurrent units (GRU) as the update function $\phi$ and are briefly explained in App. B. As an alternative to sequential models, DAGs can be encoded using Transformer-based architectures, such as PACE (Dong et al., 2022) or DAGformer (Luo et al., 2023).
@@ 4.1 Attentional GNNs
Popular attentional GNNs include GAT (Veličković et al., 2018), GATv2 (Brody et al., 2022), and TransformerConv (TC)  [^2] Please note that TransformerConv is not a Graph Transformer with a global attention module (such as GraphGPS (Rampášek et al., 2022)) but a message-passing GNN with a local attention mechanism adopted from the vanilla Transformer (Vaswani et al., 2017). It was originally introduced under the name Graph Transformer as part of the UniMP model (Shi et al., 2021). (Shi et al., 2021), which mainly differ in the choice of the scoring function $e$ (see App. LABEL:app:scoring). The graph attention mechanism is visualized in Fig. 2a.
@@ 5.2 Flow-Attention on DAGs
If all nodes in a graph have only one outgoing edge, e.g., the graph is a rooted tree, then $\beta_{ij}=1~\forall~i,~j$. In this case, the flow attention mechanism degenerates to a sum aggregation, which can be maximally expressive for the right choice of $\phi$ and $f$ (Xu et al., 2019).
@@ 5.3 FlowDAGNN
We propose a flow-attentional version of the attention-based DAGNN (Thost & Chen, 2021). A naive approach would be to simply replace the attention weights $\alpha_{ij}$ with flow attention weights $\beta_{ij}$. Due to the sequential nature of directed acyclic GNNs, the computation of the flow attention weight $\beta_{ij}$ in a DAG only depends on the node $i$ and all its ancestors. However, in contrast to standard attention weights, flow attention weights are forward-directed, which means that they should also be conditioned on all descendants of the node $i$. Analogously, the electrical current splitting up from one node into multiple branches depends on the whole branch and not only on the first node in each branch. In App. LABEL:app:example, we give a simple example of an electronic circuit illustrating this situation.
Hence, we construct a FlowDAGNN layer from two sublayers (see Fig. 2c). In the first sublayer (we call it the reverse pass), we invert all edges of the DAG $\mathcal{D}$ and apply a standard DAGNN layer to the reverse DAG $\tilde{\mathcal{D}}$. This is equivalent to performing the aggregation over all successor nodes in the original DAG $\mathcal{D}$ instead of over all predecessors: In flow attention, attention scores are normalized across outgoing neighbors, guiding a model to learn how to distribute a node’s message among them. This is especially useful on flow graphs governed by Kirchhoff’s first law, such as power grids, electronic circuits, or traffic networks. In contrast, other types of graph datasets, such as citation graphs or social networks, do not have an inherent structure that requires an information distribution across outgoing nodes. For example, there may be nodes that broadcast information to many adjacent nodes, e.g., influencers in social networks (Chen et al., 2009). Furthermore, as pointed out in the Introduction, some non-isomorphic graph structures might be equivalent in the context of informational graphs (e.g., computational graphs) and therefore should be mapped to the same representation, a consistency not guaranteed by a flow-attentional GNN due to Theorem 5.4. We therefore expect FlowGNNs to be less effective on these types of graph datasets. On the other hand, the flow-attention mechanism has the advantage over standard attention that it can distinguish between node multisets with the same distribution (Lemma 5.2). This could lead to performance improvements on graphs where certain node features occur repeatedly, which is often the case in flow graph applications (e.g., multiple identical components in electronic circuits), but can also be the case outside of the flow graph context (e.g., multiple identical atoms in molecules). In this study, we focus on evaluating the effectiveness of the flow attention mechanism on flow graph datasets, including power grids and electronic circuits. However, additional experimental results on standard graph datasets can be found in App. G.3.
@@ 6.1 Graph Classification on Undirected Flow Graphs
Dataset. We use the publicly available power grid data from the PowerGraph benchmark dataset (Varbella et al., 2024), which encompasses the IEEE24, IEEE39, IEEE118, and UK transmission systems. The graphs contained in these datasets are undirected and cyclic and represent test power systems mirroring real-world power grids. The test systems differ in scale and topology, covering various relevant parameters. Further details can be found in App. LABEL:app:data.
@@
Models and Baselines. We take three widely used attention-based GNNs (GAT, GATv2, and TransformerConv (TC)) and compare them against the corresponding flow-attentional variants FlowGAT, FlowGATv2, and FlowTC. Additionally, we compare against three popular non-attentional GNN baselines (GCN, GraphSAGE, and GIN). For each model, we perform a hyperparameter optimization by varying the number of message-passing layers (1, 2, 3) and the hidden dimension (8, 16, 32). Between subsequent message-passing layers, we apply the ReLU activation function followed by a dropout of 10%. To obtain graph embeddings from the node embeddings, we apply a global maximum pooling operator. For the final prediction, we use a single linear layer or a two-layer perceptron with a LeakyReLU activation, depending on which type of prediction layer was used for the corresponding model in the original PowerGraph benchmark. The message-passing GNNs compared in this experiment are important components of many recent Graph Transformer models Shehzad et al. (2024), which combine local message-passing with global attention. In App. LABEL:app:transformers, we report additional experimental results with recent state-of-the-art Graph Transformer models: SAT (Chen et al., 2022), GraphGPS (Rampášek et al., 2022), and Exphormer (Shirzad et al., 2023). We will investigate the integration of flow-attentional GNNs into Graph Transformer models in future work.
Experimental Setting. We stick closely to the original benchmark setting in Varbella et al. (2024) by splitting the datasets into train/validation/test with ratios 85/5/10% and using the Adam optimizer (Kingma, 2014) with an initial learning rate of $10^{-3}$ as well as a scheduler that reduces the learning rate by a factor of five if the validation accuracy plateaus for ten epochs. The negative log-likelihood is used as the loss function, and balanced accuracy (Brodersen et al., 2010) is used as the primary evaluation metric due to the strong class imbalance (see App. LABEL:app:data). We train all models with a batch size of 16 for a maximum number of 500 epochs and apply early stopping with a patience of 20 epochs. Each training run is repeated five times with different random seeds.
@@
Results. The balanced accuracies and macro-F1 scores on the test set are reported for each model on each of the four test systems in Tab. 1. We only report the results for the best model architecture from the hyperparameter optimization. Thereby, we noticed that the accuracy mostly improves with more message-passing layers, which has already been observed for power grid data in Ringsquandl et al. (2021). The FlowGNNs perform better than their corresponding standard GNN version in most cases. FlowGAT shows a higher balanced accuracy compared to GAT for the test systems IEEE24, IEEE39, and IEEE118, and a comparable performance on the UK test system. In the case of GATv2, the FlowGNN version even outperforms its standard counterpart on all test systems. FlowTC performs better than TC on all test systems except for IEEE118, where it shows a comparable performance. On all test systems, the best-performing model among the tested ones is a flow-attentional GNN. Overall, these results indicate that the flow attention mechanism, which is the only applied change to the corresponding baselines, can enhance the performance of attention-based GNNs on undirected flow graph data.
@@ 6.2 Graph Regression on Directed Acyclic Flow Graphs
Dataset. We utilize the Ckt-Bench101 dataset from the publicly available Open Circuit Benchmark (OCB) (Dong et al., 2023), which was developed to evaluate methods for electronic design automation. The dataset contains 10,000 operational amplifiers (Op-Amps) represented as DAGs and provides circuit specifications (e.g., gain and bandwidth) obtained from simulations. Details can be found in App. LABEL:app:data.
@@
Models and Baselines. We compare our proposed model, FlowDAGNN, against widely used baseline models from the literature, including GNN- and Transformer-based models tailored to DAGs: D-VAE, DAGNN, DAGformer (building upon the Structure-Aware Transformer (SAT, Chen et al. (2022))), and PACE. Thereby, we use the default parameters from Dong et al. (2023) where applicable, and from the original publications elsewhere, as well as the model-specific readout layers. For FlowDAGNN, we use two layers as described in Sec. 5.3 (each comprising one reverse and one forward pass) and adopt all other model parameters from DAGNN. The final prediction is done using a two-layer perceptron with a ReLU activation in between. Right before these final layers, we apply a dropout of 50% for regularization.
Experimental setting. We split the dataset into train/validation/test with ratios 80/10/10% and select the same test set as in Dong et al. (2023). Furthermore, we use the AdamW optimizer (Loshchilov, 2017) with an initial learning rate of $10^{-4}$ and train each model using the mean squared error (MSE) as the loss function with a batch size of 64 for a maximum of 500 epochs, but apply early stopping with a patience of 20 epochs. Each training run is repeated ten times with different random seeds.

### v3

@@ 1 Introduction
Graph neural networks (GNNs) (scarselli2008graph; kipf2017semi) have emerged as a powerful framework that extends the scope of deep learning from Euclidean to graph-structured data, which is prevalent across many real-world domains, such as social networks (fan2019graph), recommender systems (wu2022graph), materials science (reiser2022graph), or epidemiology (liu2024review). Especially attention-based GNNs have become increasingly popular due to their ability to select relevant features adaptively (sun2023attention). As graph data becomes increasingly common, advancing GNN architectures is crucial for improving performance in tasks such as node classification (hamilton2017inductive), graph regression (gilmer2017neural), or link prediction (zhang2018link).
In many important applications of GNNs, graphs are naturally associated with a flow of physical resources, such as electrical current in electronic circuits (sanchez2023comprehensive) or power grids (liao2021review), traffic in transportation networks (jiang2022graph), water in river networks (sun2021explore), or raw materials and goods in supply chains (kosasih2022machine). All nodes in these resource flow graphs, except for source and target nodes, are subject to Kirchhoff’s first law, which states that the sum of all incoming and outgoing flows must be zero, reflecting the conservation of resources. By contrast, informational graphs—such as computation graphs, social networks, or citation networks—are not associated with any physical flow but rather represent relationships or information transfer. Information can be arbitrarily duplicated in these graphs, unlike in flow graphs, where such duplication would violate the conservation law.
@@
Figure 1: Two non-isomorphic graphs that are equivalent as informational graphs, but different as resource flow graphs. a The two different directed graph structures represent the same computation (example adapted from zhang2019dvae). b The same graph structures as above represent different electronic circuits.
@@ 2 Related Work
In recent years, many new GNN models have been specifically designed for different graph types (thomas2023graph). Despite their fundamental differences, informational graphs and flow graphs are mainly treated by the same basic message-passing layers, such as GraphSAGE (hamilton2017inductive), GAT (velivckovic2018graph), or GIN (xu2019powerful). In these models, messages exchanged between neighboring nodes do not depend on the number of recipients. Instead, the information is arbitrarily duplicated and passed to all neighbors. GCN (kipf2017semi) applies a symmetric neighborhood normalization but exhibits limited expressivity and performance. Attention-based GNNs, such as GAT, GATv2 (brody2021attentive), or UniMP (shi2020masked), adaptively weight neighboring node features, leading to improved representation learning. However, the attention weights are obtained through normalization across incoming neighbors, allowing for arbitrary message duplication. Our approach normalizes across outgoing neighbors instead, which better captures the physical flow of a graph while preserving the benefits of graph attention.
@@
Many flow graphs, including the example graphs in Fig. 1, can be naturally expressed as DAGs, e.g., operational amplifiers (dong2023cktgnn) or material flow networks (perera2018topological). Before GNNs were introduced, recursive neural networks were applied to DAGs (sperduti1997supervised; frasconi1998general), and contextual recursive cascade correlation was proposed to overcome limitations in expressivity (hammer2005universal). However, these early works lacked the advantages of modern GNNs, which have been extended to DAGs in recent years (zhang2019dvae; thost2021directed). In directed acyclic GNNs, nodes are typically updated sequentially following the partial order of the DAG, and the final target node representations are used as the graph embedding. Although these sequential models outperform undirected GNNs on DAG datasets, they still aggregate node neighborhood information similarly. This means they only encode the computation tree of the output nodes and not the exact structure of the DAG, resulting in limited expressivity.
@@
A possible approach to overcome the problem of indistinguishable flow graphs is to use node indices or random features as input node features (loukas2019graph; sato2021random), which enables the model to uniquely identify each node. However, the resulting GNN model is no longer permutation invariant, which reduces its generalization capability. Similar problems arise for Transformer-based models (vaswani2017attention) such as PACE (dong2022pace), which incorporate the relational inductive bias (battaglia2018relational) via positional encodings. A different strategy would be to introduce Kirchhoff’s first law through an additional physics-informed loss term (donon2020neural), which considerably increases the training complexity and is only useful if the target variable is the resource flow itself. Our approach enhances the expressivity of attention-based GNNs on flow graphs while preserving permutation invariance and computational efficiency.
@@ 3 Preliminaries
Node Multiset. For each node in a graph, the feature vectors of a set of incoming nodes can be represented as a multiset (xu2019powerful). A multiset is a pair $(S,m)$, where $S$ is a set of distinct elements (the node features) and $m:S\rightarrow\mathbb{N}$ is the multiplicity of each element. We call two multisets $X_{1}=(S,m_{1}),~X_{2}=(S,m_{2})$ equally distributed if $m_{2}=k\cdot m_{1}$ with $k\in\mathbb{N}_{\geq 1}$.
@@
The most widely used directed acyclic GNNs are D-VAE (zhang2019dvae) and DAGNN (thost2021directed), the latter of which uses standard attention for aggregation. Both models utilize gated recurrent units (GRU) as the update function $\phi$ and are briefly explained in App. B. As an alternative to sequential models, DAGs can be encoded using Transformer-based architectures, such as PACE (dong2022pace) or DAGformer (luo2023transformers).
@@ 4.1 Attentional GNNs
Popular attentional GNNs include GAT (velivckovic2018graph), GATv2 (brody2021attentive), and TransformerConv (TC)  [^2] Please note that TransformerConv is not a Graph Transformer with a global attention module (such as GraphGPS (rampavsek2022recipe)) but a message-passing GNN with a local attention mechanism adopted from the vanilla Transformer (vaswani2017attention). It was originally introduced under the name Graph Transformer as part of the UniMP model (shi2020masked). (shi2020masked), which mainly differ in the choice of the scoring function $e$ (see App. LABEL:app:scoring). The graph attention mechanism is visualized in Fig. 2a.
@@ 5.2 Flow-Attention on DAGs
If all nodes in a graph have only one outgoing edge, e.g., the graph is a rooted tree, then $\beta_{ij}=1~\forall~i,~j$. In this case, the flow attention mechanism degenerates to a sum aggregation, which can be maximally expressive for the right choice of $\phi$ and $f$ (xu2019powerful).
@@ 5.3 FlowDAGNN
We propose a flow-attentional version of the attention-based DAGNN (thost2021directed). A naive approach would be to simply replace the attention weights $\alpha_{ij}$ with flow attention weights $\beta_{ij}$. Due to the sequential nature of directed acyclic GNNs, the computation of the flow attention weight $\beta_{ij}$ in a DAG only depends on the node $i$ and all its ancestors. However, in contrast to standard attention weights, flow attention weights are forward-directed, which means that they should also be conditioned on all descendants of the node $i$. Analogously, the electrical current splitting up from one node into multiple branches depends on the whole branch and not only on the first node in each branch. In App. LABEL:app:example, we give a simple example of an electronic circuit illustrating this situation.
Hence, we construct a FlowDAGNN layer from two sublayers (see Fig. 2c). In the first sublayer (we call it the reverse pass), we invert all edges of the DAG $\mathcal{D}$ and apply a standard DAGNN layer to the reverse DAG $\tilde{\mathcal{D}}$. This is equivalent to performing the aggregation over all successor nodes in the original DAG $\mathcal{D}$ instead of over all predecessors: In flow attention, attention scores are normalized across outgoing neighbors, guiding a model to learn how to distribute a node’s message among them. This is especially useful on flow graphs governed by Kirchhoff’s first law, such as power grids, electronic circuits, or traffic networks. In contrast, other types of graph datasets, such as citation graphs or social networks, do not have an inherent structure that requires an information distribution across outgoing nodes. For example, there may be nodes that broadcast information to many adjacent nodes, e.g., influencers in social networks (chen2009efficient). Furthermore, as pointed out in the Introduction, some non-isomorphic graph structures might be equivalent in the context of informational graphs (e.g., computational graphs) and therefore should be mapped to the same representation, a consistency not guaranteed by a flow-attentional GNN due to Theorem 5.4. We therefore expect FlowGNNs to be less effective on these types of graph datasets. On the other hand, the flow-attention mechanism has the advantage over standard attention that it can distinguish between node multisets with the same distribution (Lemma 5.2). This could lead to performance improvements on graphs where certain node features occur repeatedly, which is often the case in flow graph applications (e.g., multiple identical components in electronic circuits), but can also be the case outside of the flow graph context (e.g., multiple identical atoms in molecules). In this study, we focus on evaluating the effectiveness of the flow attention mechanism on flow graph datasets, including power grids and electronic circuits. However, additional experimental results on standard graph datasets can be found in App. G.3.
@@ 6.1 Graph Classification on Undirected Flow Graphs
Dataset. We use the publicly available power grid data from the PowerGraph benchmark dataset (varbella2024powergraph), which encompasses the IEEE24, IEEE39, IEEE118, and UK transmission systems. The graphs contained in these datasets are undirected and cyclic and represent test power systems mirroring real-world power grids. The test systems differ in scale and topology, covering various relevant parameters. Further details can be found in App. LABEL:app:data.
@@
Models and Baselines. We take three widely used attention-based GNNs (GAT, GATv2, and TransformerConv (TC)) and compare them against the corresponding flow-attentional variants FlowGAT, FlowGATv2, and FlowTC. Additionally, we compare against three popular non-attentional GNN baselines (GCN, GraphSAGE, and GIN). For each model, we perform a hyperparameter optimization by varying the number of message-passing layers (1, 2, 3) and the hidden dimension (8, 16, 32). Between subsequent message-passing layers, we apply the ReLU activation function followed by a dropout of 10%. To obtain graph embeddings from the node embeddings, we apply a global maximum pooling operator. For the final prediction, we use a single linear layer or a two-layer perceptron with a LeakyReLU activation, depending on which type of prediction layer was used for the corresponding model in the original PowerGraph benchmark. The message-passing GNNs compared in this experiment are important components of many recent Graph Transformer models shehzad2024graph, which combine local message-passing with global attention. In App. LABEL:app:transformers, we report additional experimental results with recent state-of-the-art Graph Transformer models: SAT (chen2022structure), GraphGPS (rampavsek2022recipe), and Exphormer (shirzad2023exphormer). We will investigate the integration of flow-attentional GNNs into Graph Transformer models in future work.
Experimental Setting. We stick closely to the original benchmark setting in varbella2024powergraph by splitting the datasets into train/validation/test with ratios 85/5/10% and using the Adam optimizer (kingma2014adam) with an initial learning rate of $10^{-3}$ as well as a scheduler that reduces the learning rate by a factor of five if the validation accuracy plateaus for ten epochs. The negative log-likelihood is used as the loss function, and balanced accuracy (brodersen2010balanced) is used as the primary evaluation metric due to the strong class imbalance (see App. LABEL:app:data). We train all models with a batch size of 16 for a maximum number of 500 epochs and apply early stopping with a patience of 20 epochs. Each training run is repeated five times with different random seeds.
@@
Results. The balanced accuracies and macro-F1 scores on the test set are reported for each model on each of the four test systems in Tab. 1. We only report the results for the best model architecture from the hyperparameter optimization. Thereby, we noticed that the accuracy mostly improves with more message-passing layers, which has already been observed for power grid data in ringsquandl2021power. The FlowGNNs perform better than their corresponding standard GNN version in most cases. FlowGAT shows a higher balanced accuracy compared to GAT for the test systems IEEE24, IEEE39, and IEEE118, and a comparable performance on the UK test system. In the case of GATv2, the FlowGNN version even outperforms its standard counterpart on all test systems. FlowTC performs better than TC on all test systems except for IEEE118, where it shows a comparable performance. On all test systems, the best-performing model among the tested ones is a flow-attentional GNN. Overall, these results indicate that the flow attention mechanism, which is the only applied change to the corresponding baselines, can enhance the performance of attention-based GNNs on undirected flow graph data.
@@ 6.2 Graph Regression on Directed Acyclic Flow Graphs
Dataset. We utilize the Ckt-Bench101 dataset from the publicly available Open Circuit Benchmark (OCB) (dong2023cktgnn), which was developed to evaluate methods for electronic design automation. The dataset contains 10,000 operational amplifiers (Op-Amps) represented as DAGs and provides circuit specifications (e.g., gain and bandwidth) obtained from simulations. Details can be found in App. LABEL:app:data.
@@
Models and Baselines. We compare our proposed model, FlowDAGNN, against widely used baseline models from the literature, including GNN- and Transformer-based models tailored to DAGs: D-VAE, DAGNN, DAGformer (building upon the Structure-Aware Transformer (SAT, chen2022structure)), and PACE. Thereby, we use the default parameters from dong2023cktgnn where applicable, and from the original publications elsewhere, as well as the model-specific readout layers. For FlowDAGNN, we use two layers as described in Sec. 5.3 (each comprising one reverse and one forward pass) and adopt all other model parameters from DAGNN. The final prediction is done using a two-layer perceptron with a ReLU activation in between. Right before these final layers, we apply a dropout of 50% for regularization.
Experimental setting. We split the dataset into train/validation/test with ratios 80/10/10% and select the same test set as in dong2023cktgnn. Furthermore, we use the AdamW optimizer (loshchilov2017decoupled) with an initial learning rate of $10^{-4}$ and train each model using the mean squared error (MSE) as the loss function with a batch size of 64 for a maximum of 500 epochs, but apply early stopping with a patience of 20 epochs. Each training run is repeated ten times with different random seeds.